A model for a company's revenue from selling a software package is
r = -2.5p (squared) + 500p, where p is the price in dollars of the software. What price will maximize revenue? Find the maximum revenue.

Answer by stanbon(75887)

A model for a company's revenue from selling a software package is
r = -2.5p (squared) + 500p, where p is the price in dollars of the software. What price will maximize revenue? Find the maximum revenue.
-----------------------
Binomial with a = -2.5 and b = 500
---
Max occurs at x = -b/2a = -500/(-5) = 100
---------------------
Max is r(100) = -2.5(100^2) + 500(100 = 25000
===================================================
cheers,
Stan H.